OpenStudy (anonymous):

You have $1200 to buy fencing, which cost $10 per horizontal foot. The fence will surround a rectangular field and split down the middle. What are the dimensions that will enclose the maximum possible area?

OpenStudy (anonymous):

It seems that you can buy a maximum of 120 feet of fencing with your budget. Your rectangle will have a length of a and a width of b. The rectangle itself will have a perimeter of 2a+2b. You also have a split down the middle, another length of a. The total amount of fencing you need is 3a+2b. The area of your field is a*b. So... 3a+2b=120 a*b need to be maximized. 2b=120-3a b=60-1.5a a(60-1.5a) 60a-1.5a^2 maximize by setting the derivative equal to 0 60-3a=0 60=3a a=20 3(20)+2b=120 2b=60 b=30 The dimensions are 20x30.